# Reshaping Neural Representation via Associative, Presynaptic Short-Term Plasticity

### v1

@@ D.2 Linear response to sinusoidal modulation
$\left.\frac{\mathrm{d}C_{*}^{U}}{\mathrm{d}\nu}\right|_{\nu=\nu_{0}}=\frac{1-\tau_{d}\nu_{0}U}{\bigl(1+\tau_{d}\nu_{0}U\bigr)^{3}}=\frac{2-\tau_{d}\kappa}{(\tau_{d}\kappa)^{3}}=\lim_{\omega\to 0}H_{C}^{U}(\omega),$ (D.32)
@@
$\lim_{\omega\to\infty}H_{C}^{U}(\omega)=\frac{1}{(\tau_{d}\kappa)^{2}}=\frac{1}{\bigl(1+\tau_{d}\nu_{0}U\bigr)^{2}}=f_{*}^{U},$ (D.33)
@@
D.2.5 Dependence on oscillation frequency and STP operating point.
@@
$x:=\frac{\omega}{\kappa}.$ (D.35)
@@
$\phi_{w_{0}}(\omega)=\arg H_{C}^{w_{0}}(\omega)=\arctan(\omega\tau_{d})-\arctan\!\left(\frac{\omega}{\kappa}\right)$ (D.36)
@@
$\phi_{U}(\omega)=\phi_{U}(x)=\arctan(rx)+\arctan\!\left(\frac{x}{2-r}\right)-2\arctan(x)$ (D.37)

### v2

@@ 5.4.2 Linear response to sinusoidal modulation
$\left.\frac{\mathrm{d}C_{*}^{U}}{\mathrm{d}\nu}\right|_{\nu=\nu_{0}}=\frac{1-\tau_{d}\nu_{0}U}{\bigl(1+\tau_{d}\nu_{0}U\bigr)^{3}}=\frac{2-\tau_{d}\kappa}{(\tau_{d}\kappa)^{3}}=\lim_{\omega\to 0}H_{C}^{U}(\omega),$ (S73)
@@
$\lim_{\omega\to\infty}H_{C}^{U}(\omega)=\frac{1}{(\tau_{d}\kappa)^{2}}=\frac{1}{\bigl(1+\tau_{d}\nu_{0}U\bigr)^{2}}=f_{*}^{U},$ (S74)
@@
Dependence on oscillation frequency and STP operating point.
@@
$x:=\frac{\omega}{\kappa}.$ (S76)
@@
$\phi_{w_{0}}(\omega)=\arg H_{C}^{w_{0}}(\omega)=\arctan(\omega\tau_{d})-\arctan\!\left(\frac{\omega}{\kappa}\right)$ (S77)
@@
$\phi_{U}(\omega)=\phi_{U}(x)=\arctan(rx)+\arctan\!\left(\frac{x}{2-r}\right)-2\arctan(x)$ (S78)